𝑷(𝑿 = 𝒙)=k/x+1;where x = 1,2,3


1
Expert's answer
2020-12-21T18:09:01-0500

I assume that the question asks for the value of k.

"\sum P(xi)=1"

Thus "\frac{k}{2}+ \frac{k}{3}+ \frac{k}{4}=1"

"\frac{6k+4k+3k}{12}=1"

"\frac{13k}{12}=1"

"K=\frac{12}{13}"